OpenStudy (firejay5):

Solve each system of equations by using substitution: #17 is no solution and I need the work and steps shown of how it's a no solution. 17. 6x - 3y = 12 2x = 8 - y

OpenStudy (firejay5):

@ArkGoLucky Do you have the problem/answers that went with this

OpenStudy (anonymous):

You already know how to do this don't you?

OpenStudy (firejay5):

learner

OpenStudy (anonymous):

sorry improper grammar.

OpenStudy (firejay5):

what do you mean???

OpenStudy (anonymous):

is there supposed to be a y somewhere in the second equation?

OpenStudy (firejay5):

there

OpenStudy (anonymous):

where lol

OpenStudy (firejay5):

I changed it

OpenStudy (anonymous):

okay now solve for a variable, preferably one without a coefficient. In this case, you would probably want to solve for y in the second equation.

OpenStudy (firejay5):

@ArkGoLucky Would it look like this: -y = -2x + 8 ----> y = 2x - 8

OpenStudy (anonymous):

actually, it would look like y=-2x+8 2x=8-y 2x-8=-y -2x+8=y

OpenStudy (firejay5):

show your work on how you got it @ArkGoLucky

OpenStudy (anonymous):

that's my work. that's exactly how I got it

OpenStudy (firejay5):

@ArkGoLucky do I subtract 8 to get -y

OpenStudy (anonymous):

yeah

OpenStudy (firejay5):

about

OpenStudy (firejay5):

@ArkGoLucky How is it no solution

hartnn (hartnn):

factor out 3 from first equation 6x-3y=12 what do u get ?

OpenStudy (firejay5):

I get 6 = 8 - y

OpenStudy (firejay5):

is that right

hartnn (hartnn):

nopes, 3(2x-y)=12 so 2x-y=12/3 2x-y=4 got this ?

OpenStudy (firejay5):

uh???

OpenStudy (firejay5):

I started with the bottom equation

hartnn (hartnn):

but bottom equation does not have 6, .....where did u get 6 from?

OpenStudy (firejay5):

I got 6 from, x = 3 and enter it into bottom equation

hartnn (hartnn):

but why did u put x=3 ?

OpenStudy (anonymous):

so now put y into the other equation and solve for x then you will see why there is not solution

OpenStudy (firejay5):

I got x = 3 when I substituted -2x + 8 into top equation

hartnn (hartnn):

6x-3(-2x+8)=12 like this ?

OpenStudy (firejay5):

yes

hartnn (hartnn):

6x +6x -24 =12 12x=12+24 x=3 put x=3 in 2x=8-y what do u get

OpenStudy (firejay5):

6 = 8 - y

hartnn (hartnn):

so y=?

hartnn (hartnn):

what do u subtract from 8 to get 6 ?

OpenStudy (firejay5):

???

hartnn (hartnn):

6=8-y subtract 8 from both sides......

OpenStudy (firejay5):

-2 = -y

hartnn (hartnn):

so y=?

OpenStudy (firejay5):

2

hartnn (hartnn):

so x=3,y=2 it definitely has a solution

OpenStudy (firejay5):

@ArkGoLucky How is it no solution when there's a point (3, 2)

OpenStudy (firejay5):

@hartnn The book says it's a no solution

hartnn (hartnn):

either verify the question, or verify that u are seeing correct answer